OpenStudy (anonymous):

eval the expression and write in the form a+bi . (sqrt3-sqrt-4)(sqrt6-sqrt-8)

OpenStudy (anonymous):

\[-\sqrt{2}-4i \sqrt{6} \] given \[\sqrt{-1} = i\]

OpenStudy (anonymous):

can u help with the order of multiplying